Exercise 27.8 A particle with charge -5.90 ncis moving in a uniform magnetic seld -(1.30 T)k The magnetic force on the partcle is measured to be Part B Calculate the x component of the velocity of the particle. Submit My Answers Give Up Part C Calculate the y component of the velocity of the particle. My Answers Give Up Part D Calculate the scalar product G.F m/s m/s m/s. N
media%2F8c3%2F8c341140-16f6-4b8d-9222-d0
Answer #1

F = q ( V X B)

- (3.10 x 10^ -7 ) i + (7.60 x 10^-7 ) j = -5.90 x 10^-9 ( {Vx i + Vyj} x { -1.30 k} )

(-3.10 x 10^ -7 ) i + (7.60 x 10^-7 ) j = 5.90 x 10^-9 ( {-Vx i - Vyj} x {-1.30 k} )

Taking cross product and equating,

7.60 x 10^-7 = - 5.90 x 10^-9 Vx (1.30), Vx= - 99.08 i m/s

-3.10 x 10^ -7 = 5.90 x 10^-9 Vy (1.30), Vy = -40.417 j m/s

c) v. F =- ( 99.08 i + 40.417j ) . (- 3.10 x 10^ -7 ) i + (7.60 x 10^-7 ) j= 10^-7 ( 307.148 - 307.1692) = - 2.12 x 10^ -9 m/s N

d) theta = cos^-1 (  - 2.12 x 10^ -9 / magnitude (F) magitude (v)) = cos^-1 (  - 2.12 x 10^ -9/ {107 x 8.21 x 10^-7) )

cos^-1 (- 0.2413) = 103.96 degree apprx